# A Fisher-Rao gradient flow for entropic mean-field min-max games

### v1

@@ A.2. Existence and uniqueness of the FR flow
$\int_{\mathcal{X}}\left(\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-2)},\mu_{s}^{(n-2)},x)\right)\nu_{t}^{(n)}(\mathrm{d}x)\\ =\int_{\mathcal{X}}\Bigg{(}\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)},x)\\ +\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-2)},\mu_{s}^{(n-2)},x)\Bigg{)}\nu_{t}^{(n)}(\mathrm{d}x)\\ =\int_{\mathcal{X}}\int_{\mathcal{Y}}\int_{0}^{1}\frac{\delta^{2}F}{\delta\mu\delta\nu}\left(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)}+\lambda\left(\mu_{s}^{(n-1)}-\mu_{s}^{(n-2)}\right),x,w\right)\mathrm{d}\lambda\left(\mu_{s}^{(n-1)}-\mu_{s}^{(n-2)}\right)(\mathrm{d}w)\nu_{t}^{(n)}(\mathrm{d}x)\\ +\int_{\mathcal{X}}\int_{\mathcal{X}}\int_{0}^{1}\frac{\delta^{2}F}{\delta\nu^{2}}\left(\nu_{s}^{(n-2)}+\lambda\left(\nu_{s}^{(n-1)}-\nu_{s}^{(n-2)}\right),\mu_{s}^{(n-2)},x,z\right)\mathrm{d}\lambda\left(\nu_{s}^{(n-1)}-\nu_{s}^{(n-2)}\right)(\mathrm{d}z)\nu_{t}^{(n)}(\mathrm{d}x).$
Similarly, again from (A.3) we have
@@
(A.6) $\operatorname{D_{KL}}(\nu_{t}^{(n-1)}|\nu_{t}^{(n)})=-\int_{0}^{t}\frac{\sigma^{2}}{2}e^{-\frac{\sigma^{2}}{2}(t-s)}\times\\ \times\Bigg{[}\frac{2}{\sigma^{2}}\int_{\mathcal{X}}\left(\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-2)},\mu_{s}^{(n-2)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)\right)\nu_{t}^{(n-1)}(\mathrm{d}x)\\ -\operatorname{D_{KL}}(\nu_{s}^{(n-2)}|\pi)+\operatorname{D_{KL}}(\nu_{s}^{(n-1)}|\pi)\Bigg{]}\mathrm{d}s.$
@@
$\int_{\mathcal{X}}\left(\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-2)},\mu_{s}^{(n-2)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)\right)\nu_{t}^{(n-1)}(\mathrm{d}x)\\ =-\int_{\mathcal{X}}\Bigg{(}\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)},x)\\ +\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-2)},\mu_{s}^{(n-2)},x)\Bigg{)}\nu_{t}^{(n-1)}(\mathrm{d}x)\\ =-\int_{\mathcal{X}}\int_{\mathcal{Y}}\int_{0}^{1}\frac{\delta^{2}F}{\delta\mu\delta\nu}\left(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)}+\lambda\left(\mu_{s}^{(n-1)}-\mu_{s}^{(n-2)}\right),x,w\right)\times\mathrm{d}\lambda\left(\mu_{s}^{(n-1)}-\mu_{s}^{(n-2)}\right)(\mathrm{d}w)\nu_{t}^{(n-1)}(\mathrm{d}x)\\ -\int_{\mathcal{X}}\int_{\mathcal{X}}\int_{0}^{1}\frac{\delta^{2}F}{\delta\nu^{2}}\left(\nu_{s}^{(n-2)}+\lambda\left(\nu_{s}^{(n-1)}-\nu_{s}^{(n-2)}\right),\mu_{s}^{(n-2)},x,z\right)\times\mathrm{d}\lambda\left(\nu_{s}^{(n-1)}-\nu_{s}^{(n-2)}\right)(\mathrm{d}z)\nu_{t}^{(n-1)}(\mathrm{d}x).$
Combining (A.5) and (A.6), we obtain
$\operatorname{D_{KL}}(\nu_{t}^{(n)}|\nu_{t}^{(n-1)})+\operatorname{D_{KL}}(\nu_{t}^{(n-1)}|\nu_{t}^{(n)})=-\int_{0}^{t}e^{-\frac{\sigma^{2}}{2}(t-s)}\times\Bigg{[}\\ \int_{\mathcal{X}}\int_{\mathcal{Y}}\int_{0}^{1}\frac{\delta^{2}F}{\delta\mu\delta\nu}\left(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)}+\lambda\left(\mu_{s}^{(n-1)}-\mu_{s}^{(n-2)}\right),x,w\right)\mathrm{d}\lambda\left(\mu_{s}^{(n-1)}-\mu_{s}^{(n-2)}\right)(\mathrm{d}w)\times\\ \times\left(\nu_{t}^{(n)}-\nu_{t}^{(n-1)}\right)(\mathrm{d}x)\\ +\int_{\mathcal{X}}\int_{\mathcal{X}}\int_{0}^{1}\frac{\delta^{2}F}{\delta\nu^{2}}\left(\nu_{s}^{(n-2)}+\lambda\left(\nu_{s}^{(n-1)}-\nu_{s}^{(n-2)}\right),\mu_{s}^{(n-2)},x,z\right)\mathrm{d}\lambda\left(\nu_{s}^{(n-1)}-\nu_{s}^{(n-2)}\right)(\mathrm{d}z)\times\\ \times\left(\nu_{t}^{(n)}-\nu_{t}^{(n-1)}\right)(\mathrm{d}x)\Bigg{]}\mathrm{d}s.$
@@
By the Pinsker-Csizsar inequality, $\operatorname{TV}^{2}(\nu_{t}^{(n)},\nu_{t}^{(n-1)})\leq\frac{1}{2}\operatorname{D_{KL}}(\nu_{t}^{(n)}|\nu_{t}^{(n-1)})$ and hence
@@
$\operatorname{TV}(\nu_{t}^{(n)},\nu_{t}^{(n-1)})+\operatorname{TV}(\mu_{t}^{(n)},\mu_{t}^{(n-1)})\\ \leq\frac{C_{\text{max}}}{4}\int_{0}^{t}e^{-\frac{\sigma^{2}}{2}(t-s)}\Big{(}\operatorname{TV}(\mu_{s}^{(n-1)},\mu_{s}^{(n-2)})+\operatorname{TV}(\nu_{s}^{(n-1)},\nu_{s}^{(n-2)})\Big{)}\mathrm{d}s.\\ \leq\left(\frac{C_{\text{max}}}{4}\right)^{n-1}e^{-\frac{\sigma^{2}}{2}t}\int_{0}^{t}\int_{0}^{t_{1}}\ldots\int_{0}^{t_{n-2}}e^{\frac{\sigma^{2}}{2}t_{n-1}}\Big{(}\operatorname{TV}(\nu_{t_{n-1}}^{(1)},\nu_{t_{n-1}}^{(0)})+\operatorname{TV}(\mu_{t_{n-1}}^{(1)},\mu_{t_{n-1}}^{(0)})\Big{)}\mathrm{d}t_{n-1}\ldots\mathrm{d}t_{2}\mathrm{d}t_{1}\\ \leq\left(\frac{C_{\text{max}}}{4}\right)^{n-1}e^{-\frac{\sigma^{2}}{2}t}\frac{t^{n-2}}{(n-2)!}\int_{0}^{t}e^{\frac{\sigma^{2}}{2}t_{n-1}}\Big{(}\operatorname{TV}(\nu_{t_{n-1}}^{(1)},\nu_{t_{n-1}}^{(0)})+\operatorname{TV}(\mu_{t_{n-1}}^{(1)},\mu_{t_{n-1}}^{(0)})\Big{)}\mathrm{d}t_{n-1}\\ \leq\left(\frac{C_{\text{max}}}{4}\right)^{n-1}\frac{t^{n-2}}{(n-2)!}\int_{0}^{t}\Big{(}\operatorname{TV}(\nu_{t_{n-1}}^{(1)},\nu_{t_{n-1}}^{(0)})+\operatorname{TV}(\mu_{t_{n-1}}^{(1)},\mu_{t_{n-1}}^{(0)})\Big{)}\mathrm{d}t_{n-1},$
@@
For sufficiently large $n$, the constant on the right hand side becomes less than $1$ and the proof is complete. ∎
By Lemma A.2, for any $T>0$ we obtain the existence of a pair of flows $(\nu_{t},\mu_{t})_{t\in[0,T]}$ satisfying (A.1). Moreover, for Lebesgue-almost all $t\in[0,T]$ we have
@@
(A.7) $\operatorname{D_{KL}}(\nu_{t}|\pi)\leq\liminf_{n\to\infty}\operatorname{D_{KL}}(\nu_{t}^{(n)}|\pi)\leq 2\log R_{\nu}+\frac{4}{\sigma^{2}}C_{\nu}\,,$
(A.8) $\operatorname{D_{KL}}(\mu_{t}|\rho)\leq\liminf_{n\to\infty}\operatorname{D_{KL}}(\mu_{t}^{(n)}|\rho)\leq 2\log R_{\mu}+\frac{4}{\sigma^{2}}C_{\mu}\,,$
@@
Step 2: Ratio condition (19). Using (A.3) and (A.4), we see that for any $t\in[0,T]$ we have

### v2

@@ A.2 Existence and uniqueness of the FR flow
$\int_{\mathcal{X}}\left(\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-2)},\mu_{s}^{(n-2)},x)\right)\nu_{t}^{(n)}(\mathrm{d}x)\\ =\int_{\mathcal{X}}\Bigg{(}\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)},x)+\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-2)},\mu_{s}^{(n-2)},x)\Bigg{)}\nu_{t}^{(n)}(\mathrm{d}x)\\ =\int_{\mathcal{X}}\int_{\mathcal{Y}}\int_{0}^{1}\frac{\delta^{2}F}{\delta\mu\delta\nu}\left(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)}+\lambda\left(\mu_{s}^{(n-1)}-\mu_{s}^{(n-2)}\right),x,w\right)\mathrm{d}\lambda\left(\mu_{s}^{(n-1)}-\mu_{s}^{(n-2)}\right)(\mathrm{d}w)\nu_{t}^{(n)}(\mathrm{d}x)\\ +\int_{\mathcal{X}}\int_{\mathcal{X}}\int_{0}^{1}\frac{\delta^{2}F}{\delta\nu^{2}}\left(\nu_{s}^{(n-2)}+\lambda\left(\nu_{s}^{(n-1)}-\nu_{s}^{(n-2)}\right),\mu_{s}^{(n-2)},x,z\right)\mathrm{d}\lambda\left(\nu_{s}^{(n-1)}-\nu_{s}^{(n-2)}\right)(\mathrm{d}z)\nu_{t}^{(n)}(\mathrm{d}x).$
Similarly, again from (29) we have
@@
$\operatorname{D_{KL}}(\nu_{t}^{(n-1)}|\nu_{t}^{(n)})=-\int_{0}^{t}\frac{\sigma^{2}}{2}e^{-\frac{\sigma^{2}}{2}(t-s)}\Bigg{[}\frac{2}{\sigma^{2}}\int_{\mathcal{X}}\left(\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-2)},\mu_{s}^{(n-2)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)\right)\nu_{t}^{(n-1)}(\mathrm{d}x)\\ -\operatorname{D_{KL}}(\nu_{s}^{(n-2)}|\pi)+\operatorname{D_{KL}}(\nu_{s}^{(n-1)}|\pi)\Bigg{]}\mathrm{d}s.$ (32)
@@
$\int_{\mathcal{X}}\left(\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-2)},\mu_{s}^{(n-2)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)\right)\nu_{t}^{(n-1)}(\mathrm{d}x)\\ =-\int_{\mathcal{X}}\Bigg{(}\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)},x)+\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)},x)-\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-2)},\mu_{s}^{(n-2)},x)\Bigg{)}\nu_{t}^{(n-1)}(\mathrm{d}x)\\ =-\int_{\mathcal{X}}\int_{\mathcal{Y}}\int_{0}^{1}\frac{\delta^{2}F}{\delta\mu\delta\nu}\left(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)}+\lambda\left(\mu_{s}^{(n-1)}-\mu_{s}^{(n-2)}\right),x,w\right)\times\mathrm{d}\lambda\left(\mu_{s}^{(n-1)}-\mu_{s}^{(n-2)}\right)(\mathrm{d}w)\nu_{t}^{(n-1)}(\mathrm{d}x)\\ -\int_{\mathcal{X}}\int_{\mathcal{X}}\int_{0}^{1}\frac{\delta^{2}F}{\delta\nu^{2}}\left(\nu_{s}^{(n-2)}+\lambda\left(\nu_{s}^{(n-1)}-\nu_{s}^{(n-2)}\right),\mu_{s}^{(n-2)},x,z\right)\times\mathrm{d}\lambda\left(\nu_{s}^{(n-1)}-\nu_{s}^{(n-2)}\right)(\mathrm{d}z)\nu_{t}^{(n-1)}(\mathrm{d}x).$
Combining (31) and (32), we obtain
$\operatorname{D_{KL}}(\nu_{t}^{(n)}|\nu_{t}^{(n-1)})+\operatorname{D_{KL}}(\nu_{t}^{(n-1)}|\nu_{t}^{(n)})=-\int_{0}^{t}e^{-\frac{\sigma^{2}}{2}(t-s)}\times\Bigg{[}\\ \int_{\mathcal{X}}\int_{\mathcal{Y}}\int_{0}^{1}\frac{\delta^{2}F}{\delta\mu\delta\nu}\left(\nu_{s}^{(n-1)},\mu_{s}^{(n-2)}+\lambda\left(\mu_{s}^{(n-1)}-\mu_{s}^{(n-2)}\right),x,w\right)\mathrm{d}\lambda\left(\mu_{s}^{(n-1)}-\mu_{s}^{(n-2)}\right)(\mathrm{d}w)\left(\nu_{t}^{(n)}-\nu_{t}^{(n-1)}\right)(\mathrm{d}x)\\ +\int_{\mathcal{X}}\int_{\mathcal{X}}\int_{0}^{1}\frac{\delta^{2}F}{\delta\nu^{2}}\left(\nu_{s}^{(n-2)}+\lambda\left(\nu_{s}^{(n-1)}-\nu_{s}^{(n-2)}\right),\mu_{s}^{(n-2)},x,z\right)\mathrm{d}\lambda\left(\nu_{s}^{(n-1)}-\nu_{s}^{(n-2)}\right)(\mathrm{d}z)\left(\nu_{t}^{(n)}-\nu_{t}^{(n-1)}\right)(\mathrm{d}x)\Bigg{]}\mathrm{d}s.$
@@
By the Pinsker-Csizsar inequality, $\operatorname{TV}^{2}(\nu_{t}^{(n)},\nu_{t}^{(n-1)})\leq\frac{1}{2}\operatorname{D_{KL}}(\nu_{t}^{(n)}|\nu_{t}^{(n-1)}),$ and hence
@@
$\operatorname{TV}(\nu_{t}^{(n)},\nu_{t}^{(n-1)})+\operatorname{TV}(\mu_{t}^{(n)},\mu_{t}^{(n-1)})\leq\frac{C_{\text{max}}}{4}\int_{0}^{t}e^{-\frac{\sigma^{2}}{2}(t-s)}\Big{(}\operatorname{TV}(\mu_{s}^{(n-1)},\mu_{s}^{(n-2)})+\operatorname{TV}(\nu_{s}^{(n-1)},\nu_{s}^{(n-2)})\Big{)}\mathrm{d}s.\\ \leq\left(\frac{C_{\text{max}}}{4}\right)^{n-1}e^{-\frac{\sigma^{2}}{2}t}\int_{0}^{t}\int_{0}^{t_{1}}\ldots\int_{0}^{t_{n-2}}e^{\frac{\sigma^{2}}{2}t_{n-1}}\Big{(}\operatorname{TV}(\nu_{t_{n-1}}^{(1)},\nu_{t_{n-1}}^{(0)})+\operatorname{TV}(\mu_{t_{n-1}}^{(1)},\mu_{t_{n-1}}^{(0)})\Big{)}\mathrm{d}t_{n-1}\ldots\mathrm{d}t_{2}\mathrm{d}t_{1}\\ \leq\left(\frac{C_{\text{max}}}{4}\right)^{n-1}e^{-\frac{\sigma^{2}}{2}t}\frac{t^{n-2}}{(n-2)!}\int_{0}^{t}e^{\frac{\sigma^{2}}{2}t_{n-1}}\Big{(}\operatorname{TV}(\nu_{t_{n-1}}^{(1)},\nu_{t_{n-1}}^{(0)})+\operatorname{TV}(\mu_{t_{n-1}}^{(1)},\mu_{t_{n-1}}^{(0)})\Big{)}\mathrm{d}t_{n-1}\\ \leq\left(\frac{C_{\text{max}}}{4}\right)^{n-1}\frac{t^{n-2}}{(n-2)!}\int_{0}^{t}\Big{(}\operatorname{TV}(\nu_{t_{n-1}}^{(1)},\nu_{t_{n-1}}^{(0)})+\operatorname{TV}(\mu_{t_{n-1}}^{(1)},\mu_{t_{n-1}}^{(0)})\Big{)}\mathrm{d}t_{n-1},$
@@
Using the definition of $\mathcal{TV}^{[0,T]},$ the last inequality becomes
$\mathcal{TV}^{[0,T]}\left((\nu_{t}^{(n)},\mu_{t}^{(n)})_{t\in[0,T]},(\nu_{t}^{(n-1)},\mu_{t}^{(n-1)})_{t\in[0,T]}\right)\\ \leq\left(\frac{C_{\text{max}}}{4}\right)^{n-1}\frac{T^{n-1}}{(n-1)!}\mathcal{TV}^{[0,T]}\left((\nu_{t}^{(1)},\mu_{t}^{(1)})_{t\in[0,T]},(\nu_{t}^{(0)},\mu_{t}^{(0)})_{t\in[0,T]}\right).$
By choosing $n$ sufficiently large, we conclude that $\left((\nu_{t}^{(n)},\mu_{t}^{(n)})_{t\in[0,T]}\right)_{n=0}^{\infty}$ is a Cauchy sequence. By completeness of $\left(\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]},\mathcal{TV}^{[0,T]}\right),$ the sequence admits a limit point $(\nu_{t},\mu_{t})_{t\in[0,T]}\in\left(\mathcal{P}(\mathcal{X})^{[0,T]}\times\mathcal{P}(\mathcal{Y})^{[0,T]},\mathcal{TV}^{[0,T]}\right).$
Step 2: The limit point $(\nu_{t},\mu_{t})_{t\in[0,T]}$ is a fixed point of $\phi.$ From Step 1, we obtain that for Lebesgue-almost all $t\in[0,T]$ we have
$\operatorname{TV}(\nu_{t}^{(n)},\nu_{t})\to 0,\quad\operatorname{TV}(\mu_{t}^{(n)},\mu_{t})\to 0\qquad\text{ as }n\to\infty.$
Thus, by (12) and (13), we have that for Lebesgue-almost all $t\in[0,T]$ and any fixed $(x,y)\in\mathcal{X}\times\mathcal{Y},$ the flat derivatives $\frac{\delta F}{\delta\nu}(\nu_{t}^{(n-1)},\mu_{t}^{(n-1)},x)$ and $\frac{\delta F}{\delta\mu}(\nu_{t}^{(n-1)},\mu_{t}^{(n-1)},y)$ from (29) and (30) converge to $\frac{\delta F}{\delta\nu}(\nu_{t},\mu_{t},x)$ and $\frac{\delta F}{\delta\mu}(\nu_{t},\mu_{t},y),$ respectively, as $n\to\infty.$ Using Assumption 2, we can apply the dominated convergence theorem and obtain
$\lim_{n\to\infty}\int_{0}^{t}\frac{\delta F}{\delta\nu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},x)\mathrm{d}s=\int_{0}^{t}\frac{\delta F}{\delta\nu}(\nu_{s},\mu_{s},x)\mathrm{d}s,$
$\lim_{n\to\infty}\int_{0}^{t}\frac{\delta F}{\delta\mu}(\nu_{s}^{(n-1)},\mu_{s}^{(n-1)},y)\mathrm{d}s=\int_{0}^{t}\frac{\delta F}{\delta\mu}(\nu_{s},\mu_{s},y)\mathrm{d}s,$
for Lebesgue-almost all $t\in[0,T]$ and any fixed $(x,y)\in\mathcal{X}\times\mathcal{Y}.$
Moreover, in (29) and (30), using Assumption 2, (15), Lemma A.1 and the fact that $(\nu_{t}^{(n)},\mu_{t}^{(n)})_{t\in[0,T]}$ is a flow of probability measures for all $n,$ we obtain for all $n$ and $t\in[0,T],$ and any fixed $(x,y)\in\mathcal{X}\times\mathcal{Y},$ that
$\nu_{t}^{(n)}(x)\log\frac{\nu_{t}^{(n)}(x)}{\pi(x)}\leq 3\log R_{\nu}+\frac{6}{\sigma^{2}}C_{\nu},$
$\mu_{t}^{(n)}(y)\log\frac{\mu_{t}^{(n)}(y)}{\rho(y)}\leq 3\log R_{\mu}+\frac{6}{\sigma^{2}}C_{\mu}.$
Furthermore, since $(\nu_{t}^{(n)},\mu_{t}^{(n)})_{t\in[0,T]}$ are considered to be absolutely continuous with respect to the Lebesgue measure for all $n,$ it follows that convergence in TV distance implies convergence in $L^{1}.$ Then due to the dominated convergence theorem, we deduce that $\operatorname{D_{KL}}(\nu_{t}^{(n)}|\pi)$ and $\operatorname{D_{KL}}(\mu_{t}^{(n)}|\rho)$ converge to $\operatorname{D_{KL}}(\nu_{t}|\pi)$ and $\operatorname{D_{KL}}(\mu_{t}|\rho),$ respectively, for Lebesgue-almost all $t\in[0,T],$ as $n\to\infty.$ Hence, using again Lemma A.1 and the dominated convergence theorem, we obtain
$\lim_{n\to\infty}\int_{0}^{t}\operatorname{D_{KL}}(\nu_{s}^{(n)}|\pi)\mathrm{d}s=\int_{0}^{t}\operatorname{D_{KL}}(\nu_{s}|\pi)\mathrm{d}s,$
$\lim_{n\to\infty}\int_{0}^{t}\operatorname{D_{KL}}(\mu_{s}^{(n)}|\rho)\mathrm{d}s=\int_{0}^{t}\operatorname{D_{KL}}(\mu_{s}|\rho)\mathrm{d}s,$
for Lebesgue-almost all $t\in[0,T].$
Therefore, letting $n\to\infty$ in (29) and (30) and using the fact that $x\mapsto\log x$ is continuous on $(0,\infty),$ we conclude that $(\nu_{t},\mu_{t})_{t\in[0,T]}$ is a fixed point of $\phi.$
Step 3: The fixed point $(\nu_{t},\mu_{t})_{t\in[0,T]}$ of $\phi$ is unique. Suppose, for the contrary, that $\phi$ admits two fixed points $(\nu_{t},\mu_{t})_{t\in[0,T]}$ and $(\bar{\nu}_{t},\bar{\mu}_{t})_{t\in[0,T]}$ such that $\nu_{0}=\bar{\nu}_{0}$ and $\mu_{0}=\bar{\mu}_{0}.$ Then repeating the same calculations from Step 1, we arrive at
$\operatorname{D_{KL}}(\nu_{t}|\bar{\nu}_{t})+\operatorname{D_{KL}}(\bar{\nu}_{t}|\nu_{t})\leq\max\{C_{\mu,\nu},C_{\nu,\nu}\}\operatorname{TV}(\nu_{t},\bar{\nu}_{t})\int_{0}^{t}e^{-\frac{\sigma^{2}}{2}(t-s)}\left(\operatorname{TV}(\mu_{s},\bar{\mu}_{s})+\operatorname{TV}(\nu_{s},\bar{\nu}_{s})\right)\mathrm{d}s.$
By the Pinsker-Csizsar inequality, $\operatorname{TV}^{2}(\nu_{t},\bar{\nu}_{t})\leq\frac{1}{2}\operatorname{D_{KL}}(\nu_{t}|\bar{\nu}_{t}),$ and hence
$4\operatorname{TV}^{2}(\nu_{t},\bar{\nu}_{t})\leq\max\{C_{\mu,\nu},C_{\nu,\nu}\}\operatorname{TV}(\nu_{t},\bar{\nu}_{t})\int_{0}^{t}e^{-\frac{\sigma^{2}}{2}(t-s)}\left(\operatorname{TV}(\mu_{s},\bar{\mu}_{s})+\operatorname{TV}(\nu_{s},\bar{\nu}_{s})\right)\mathrm{d}s,$
which gives
$\operatorname{TV}(\nu_{t},\bar{\nu}_{t})\leq\frac{1}{4}\max\{C_{\mu,\nu},C_{\nu,\nu}\}\int_{0}^{t}e^{-\frac{\sigma^{2}}{2}(t-s)}\left(\operatorname{TV}(\mu_{s},\bar{\mu}_{s})+\operatorname{TV}(\nu_{s},\bar{\nu}_{s})\right)\mathrm{d}s,$
An almost identical argument leads to
$\operatorname{TV}(\mu_{t},\bar{\mu}_{t})\leq\frac{1}{4}\max\{C_{\nu,\mu},C_{\mu,\mu}\}\int_{0}^{t}e^{-\frac{\sigma^{2}}{2}(t-s)}\left(\operatorname{TV}(\mu_{s},\bar{\mu}_{s})+\operatorname{TV}(\nu_{s},\bar{\nu}_{s})\right)\mathrm{d}s.$
$\operatorname{TV}(\nu_{t},\bar{\nu}_{t})+\operatorname{TV}(\mu_{t},\bar{\mu}_{t})\leq\frac{C_{\text{max}}}{4}\int_{0}^{t}e^{-\frac{\sigma^{2}}{2}(t-s)}\left(\operatorname{TV}(\mu_{s},\bar{\mu}_{s})+\operatorname{TV}(\nu_{s},\bar{\nu}_{s})\right)\mathrm{d}s.$
For each $t\in[0,T],$ denote $f(t)\coloneqq\int_{0}^{t}e^{\frac{\sigma^{2}}{2}s}\left(\operatorname{TV}(\mu_{s},\bar{\mu}_{s})+\operatorname{TV}(\nu_{s},\bar{\nu}_{s})\right)\mathrm{d}s.$ Observe that $f\geq 0$ and $f(0)=0.$ Then, by Gronwall’s lemma, we obtain
$0\leq f(t)\leq e^{\frac{C_{\text{max}}}{4}t}f(0)=0,$
and hence
$\operatorname{TV}(\nu_{t},\bar{\nu}_{t})+\operatorname{TV}(\mu_{t},\bar{\mu}_{t})=0,$
for Lebesgue-almost all $t\in[0,T],$ which implies
$\nu_{t}=\bar{\nu}_{t},\quad\mu_{t}=\bar{\mu}_{t},$
for Lebesgue-almost all $t\in[0,T].$ Therefore, the fixed point $(\nu_{t},\mu_{t})_{t\in[0,T]}$ of $\phi$ must be unique.
From Steps 1, 2 and 3, we obtain the existence and uniqueness of a pair of flows $(\nu_{t},\mu_{t})_{t\in[0,T]}$ satisfying (27) for any $T>0.$
Returning to the proof of Theorem 2.2, recall from Step 1 of Lemma A.2 that for Lebesgue-almost all $t\in[0,T]$ we have
@@
$\operatorname{D_{KL}}(\nu_{t}|\pi)\leq\liminf_{n\to\infty}\operatorname{D_{KL}}(\nu_{t}^{(n)}|\pi)\leq 2\log R_{\nu}+\frac{4}{\sigma^{2}}C_{\nu}\,,$ (33)
$\operatorname{D_{KL}}(\mu_{t}|\rho)\leq\liminf_{n\to\infty}\operatorname{D_{KL}}(\mu_{t}^{(n)}|\rho)\leq 2\log R_{\mu}+\frac{4}{\sigma^{2}}C_{\mu}\,,$ (34)
@@
Step 2: Ratio condition (21). Using (29) and (30), we see that for any $t\in[0,T]$ we have